If s= 1/n, where n is any integer, what is the greatest possible value of s+100?
F. O
G. 100
H. 100.5
J. 101
K. n/A

Answer by macston(5194)

The greatest value of s if n is an integer is when n=1: s=1.
The greatest possible value of s+100=1+100=101, ANSWER: (J.)